OpenStudy (sette):

The length of a rectangle is 3 more than twice its width. The perimeter is 48 feet. Find the width. 16 7 8 21 24

OpenStudy (sette):

2l+2w=48 l=2w+3 iGot that far.

OpenStudy (anonymous):

not sure what you mean, maybe it's this 2l + 2w = 48 2x3 + 2w = 48 6 + 2w = 48 2w = 42 w = 21

OpenStudy (anonymous):

$$ l = 2w +3 $$ $$ 2(l + w) = 48 $$ $$ 2( (2w +3) + w) = $$ $$ 6w + 6 = 48 $$ $$ w = 7 , l = 17 $$

OpenStudy (anonymous):

2L+2W=48 so L+W=24 and L=3+2W so 3+2W+W=24 =>3W=21 so W=7 and therefore L=2x7+3=17

OpenStudy (anonymous):

17 is not an option, only those above